OpenStudy (anonymous):

Find the potential solution to the equation log4(2 – x) = log4(–5x – 18).

Nnesha (nnesha):

\[\log_4(2 – x) = \log_4(–5x – 18).\] for example log_4 = b \[\huge\rm b(2 – x) = b(–5x – 18).\] how would you solve this for x ?

OpenStudy (jdoe0001):

hint: UNlog both sides

OpenStudy (anonymous):

so i should distribute the 4 into each?

Nnesha (nnesha):

4 ?? 4 is base of log both sides llog bases are same

Nnesha (nnesha):

in other words divide log_4 both side

OpenStudy (anonymous):

ohh thanks

Nnesha (nnesha):

yep then just simple algebra stuff :-D